What are the next two terms of the sequence: -8; 4; -2
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

looks like the formula for the next term in the sequence is the curreent term * minus 1/3.
-8 * -1/2 = 4
4 * -1/2 = -2
-2 * -1/2 = 1
1 * -1/2 = -1/2

Answer by ikleyn(52788)

.

Next two numbers in the sequence may repeat the given numbers,
making a cyclic sequence.

Or they may be  4,  -8,  making a reflection as in the mirror.


Nobody knows a  " correct "  answer.

Why ? - Because the notion itself  " a correct answer "  is not applicable to such questions/problems.


Questions of this kind are specially intended for persons,
who have nothing to do, languish from idleness and want very much
to kill their time for nothing, without any visible benefits/reasons.


You may kill your own time as much as you want - it is your inalienable right.
But to involve others in such a deal is  OUT  of  THE  RULES  of a good style.


Not a Math problem - a typical gibberish with no sense.